OpenStudy (anonymous):

Simplify the expression: 5(3z^3-6)+15(2-z^3)

Nnesha (nnesha):

first distribute bracket by 5 and other bracket by 15

Nnesha (nnesha):

3z^3 multiply by 5 = ?? -6 multiply 5= ??? 2 multiply 15 = ?? -z^3 multiply by 15 =???

OpenStudy (xxdragoonxx):

5(3z^3-6)+15(2-z^3) 15(-z^3+2)+5(3z^3-6) = -15z^3+5(3z^3-6)+30 = 15z^3-15z^3-30+30 = 0